生物物理化学特論 2 単位 (選択)

Advanced Biophysical Chemistry  松木均 · 教授/ 環境創生工学専攻生命テクノサイエンスコース生物機能工学講座

【授業目的】生体分子の溶解した溶液の取り扱い方およびその取り扱いを用いた 生体分子集合体と生理活性物質の間に働く相互作用の解析方法を講述する． 【授業概要】本講義の前半部分では，生体分子が作用発現する場である溶液につ いて巨視的，微視的および電気化学的側面について講述する．後半部分では， 生体分子が形成する分子集合体と生理活性物質の間に働く相互作用として脂 質二分子膜中へのリガンド分配を取り上げる．上述した内容を踏まえ，過去か ら現在までの麻酔作用の分子機構について解説する．さらに相互作用の工学 的応用 (ドラッグデリバリーシステムなど) についても述べる．本科目は，工 業に関する科目である． 【授業形式】講義 【キーワード】溶液，分子集合体，リガンド，相互作用，麻酔機構 【先行科目】『物理化学 2』 (1.0)，『生物物理化学 2』 (1.0) 【関連科目】『生体熱力学』 (0.5) 【履修要件】学部教育における物理化学および生物物理化学を理解していること． 【履修上の注意】 [注意] 【到達目標】 1. 溶液の巨視的，微視的および電気化学的取り扱いを理解する． 2. 生体高分子中へのリガンドの分配様式および麻酔分子機構を理解する． 【授業計画】 1. 生体関連物質の溶液化学 (1) 溶液の熱力学 1: 基礎熱力学関係式 2. 生体関連物質の溶液化学 (2) 溶液の熱力学 2: 相図の理論 3. 生体関連物質の溶液化学 (3) 溶液の熱力学 3: 気体，液体および固体との平衡 4. 生体関連物質の溶液化学 (4) 溶液の熱力学 4: 界面現象 5. 生体関連物質の溶液化学 (5) 溶液の統計熱力学 1: 基礎統計力学関係式 6. 生体関連物質の溶液化学 (6) 溶液の統計熱力学 2: 理想溶液と正則溶液 7. 生体関連物質の溶液化学 (7) 溶液の電気化学 1: イオン平衡 8. 生体関連物質の溶液化学 (8) 溶液の電気化学 2: 電極反応 9. 生体高分子と生理活性物質の相互作用 (1) 脂質膜中へのリガンド分配 1: 非 特異的結合 10. 生体高分子と生理活性物質の相互作用 (2) 脂質膜中へのリガンド分配 2: 特 異的結合 11. 生体高分子と生理活性物質の相互作用 (3) リガンド解離平衡の影響 12. 生体高分子と生理活性物質の相互作用 (4) 麻酔機構 1:概論

13. 生体高分子と生理活性物質の相互作用 (5) 麻酔機構 2:脂質膜説

14. 生体高分子と生理活性物質の相互作用 (6) 麻酔機構 3:タンパク説

15. 生体高分子と生理活性物質の相互作用 (7) 局所麻酔薬の作用機構

16. 総括，質疑，レポート作成

【成績評価基準】出席率 80%以上で，課題レポート (100%) で評価する．【教科書】授業中に適宜資料を配布する. 【参考書】

⋄ 本村欣士著「溶液化学」朝倉書店

⋄ J. S. Rawlinson and F. L. Swinton ”Liquids and Liquid Mixtures” Butterworths Scientific

【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216738 【対象学生】開講コース学生のみ履修可能【連絡先】

⇒ 松木 (化生棟 607, 088-656-7513, matsuki@bio.tokushima-u.ac.jp) ���� (オ

フィスアワー: 金曜日 16:20-17:50) 【備考】

⋄ 授業を受ける際には， 2 時間の授業時間毎に 2 時間の予習と 2 時間の復習をしたうえで授業を受けることが，授業の理解と単位取得のために必要である．

⋄ 到達目標 1 は授業計画 1-8 に，到達目標 2 は授業計画 9-15 に関係する．

Advanced Biophysical Chemistry 2 units (selection)

Hitoshi Matsuki · Professor / Biological Functions, Biological Science and Technology, Earth and Life Environmental Engineering

Target⟩ Treatments of solutions containing biomolecules and analytical methods for interactions between biomolecular aggregates and physiologically active substances by means of the treatments are described.

Outline⟩ The former part of this lecture describes solutions in which biomolecules reveal their actions from the macroscopic, microscopic and electrochemical viewpoints. The latter part deals with ligand partitioning into lipid bilayer membranes as examples of interactions between molecular aggregates formed by biomolecules and physiologically active substances. The molecular mechanisms of anesthesia until now are reviewed on the basis of the above subjects. Further industrial application of the interactions such as drug delivery systems is also explained. This lecture deals with industrial subjects.

Style⟩ Lecture �������⟩ solution, molecular aggregate, ligand, interaction, mechanism of

anesthesia Fundamental Lecture⟩ “Physical Chemistry 2”(1.0), “Biophysical Chemistry 2”

(1.0)

Relational Lecture⟩ “Biochemical Thermodynamics”(0.5) Requirement⟩ Students are required to have a good understanding of undergraduate-

level physical chemistry and biophysical chemistry and related subjects. Notice⟩ [注意]

Goal⟩

1. To understand the macroscopic, microscopic and electrochemical treatments of solutions.

2. To understand the partitioning modes of ligands into biomacromolecules and

molecular mechanisms of anesthesia.

Schedule⟩

1. Solution chemistry of bio-related substances (1) thermodynamics of solutions 1: basic equations of thermodynamics

2. Solution chemistry of bio-related substances (2) thermodynamics of solutions 2: theory of phase diagrams

3. Solution chemistry of bio-related substances (3) thermodynamics of solutions 3: equilibrium of solutions with gases, liquids and solids 4. Solution chemistry of bio-related substances(4) thermodynamics of solutions

4: interfacial phenomena

5. Solution chemistry of bio-related substances (5) statistical thermodynamics of solutions 1: basic equations of statistical mechanics

6. Solution chemistry of bio-related substances (6) statistical thermodynamics of solutions 2: ideal and regular solutions

7. Solution chemistry of bio-related substances (7) electrochemistry of solutions 1: ionic equilibrium

8. Solution chemistry of bio-related substances (8) electrochemistry of solutions 2: electrode reaction

9. Interactions between biomacromolecules and physiologically active substances (1) ligand partitioning into lipid membranes 1: nonspecific binding

10. Interactions between biomacromolecules and physiologically active substances (2) ligand partitioning into lipid membranes 2: specific binding

11. Interactions between biomacromolecules and physiologically active substances (3) effect of dissociation equilibrium of ligand

12. Interactions between biomacromolecules and physiologically active substances (4) mechanisms of anesthesia 1: introduction

13. Interactions between biomacromolecules and physiologically active substances (5) mechanisms of anesthesia 2: lipid membrane theories

14. Interactions between biomacromolecules and physiologically active substances (6) mechanisms of anesthesia 3: protein receptor theories

15. Interactions between biomacromolecules and physiologically active substances (7) action mechanisms of local anesthetics

16. Summary, inquiry and report preparations Evaluation Criteria⟩ More than 80% percentage of attendance and reports (100%). Textbook⟩ To be distributed materials adequately in the class.

Reference⟩ ⋄ 本村欣士著「溶液化学」朝倉書店 ⋄ J. S. Rawlinson and F. L. Swinton ”Liquids and Liquid Mixtures” Butterworths Scientific

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216738

Student⟩ Able to be taken by only specified class(es)

Contact⟩ ⇒ Matsuki (G607, +81-88-656-7513, matsuki@bio.tokushima-u.ac.jp) ����

(Office Hour: Friday 16:20-17:50) Note⟩

⋄ When you take this class, it is necessary to do preparation for 2h and review for 2h everry 2h class for your understanding and taking credit. ⋄ Goal 1 is related to schedules 1-8 and goal 2 is related to schedules 9-15.